#10. Use spherical coordinates to evaluate to j** 122 de drody.
Answer #1

So E √4-x=yz 1272 dz dady -2.0. Now ocz<54-xr-ya Osv=14-42 . -2 <Y £2 Now z = √ 4-x%-y2 :> z=4-27 yr => x+yY+z=24 in sphericaso, & varies from oto 7/2 and o varies from oto T. Therefore a ,2714-42 12 22 dzdx dy -20 roo 17, T 2.0 12 (89 cos? 0.59 singTherefore - 3월 - g] - 38 151 + ] 384T 53. 지 = 804248